OpenStudy (anonymous):

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and train b is traveling at 96 miles per hour. Train A passes a station at 1:10 a.m. if train B passes the same station at 1:25 a.m., at what time will train b catch up to train A.

OpenStudy (anonymous):

I love mechanics. Find out how far the first train has travelled by 1:25. The second train has to 'catch up' this distance, and it is catching up at a speed of 96-80 = 16mph.

OpenStudy (anonymous):

how do i find that out??

OpenStudy (anonymous):

Between 1:10 and 1:25, 15 minutes passes - 1/4 hours distance = speed x time -> train A travels 80 x (1/4) = 20miles in this time. From 1:25, train B catches up at 96 - 80 = 16 mph. It needs to catch up 20 miles at 16mph relative speed, so the time is (20/16) hours after 1:25

OpenStudy (anonymous):

that is so confusing?!?!

OpenStudy (anonymous):

Just do it the old fashioned way if my mechanics is too elegantly performed for you.

OpenStudy (anonymous):

lol i dont know the old fashioned way either

OpenStudy (anonymous):

would it be 3:10 am ??